# Hire When You Need to: Gradual Participant Recruitment for Auction-based Federated Learning

### v1

@@ 3. System Model of GPS-AFL
Step 4: FL Model Training. The optimization of the global loss function is achieved by minimizing the weighted mean of the participants’ local loss functions (McMahan et al., 2017), denoted as $\ell_{n}(\omega_{n}(t))$. As such, each selected participant train its local model, $\omega_{n}(t)$, for a single time step using its own local dataset and compute resources, in attempt to minimize $\ell_{n}(\omega_{n}(t))$ within the allocated time. The actual time allocated for training will be shorter than the actual time step $t$ as time has to be allocated for participants to upload their updated model weights. We use $t_{t-r}$ to denote the actual time period used for local model training, where $0\leqslant r\leqslant t$ is the amount of communication time for participants to upload their model weights.
Step 5: FL Model Aggregation. Participants are required to submit their local model updates to the FL server $f$ at the end of $t_{t-r}$. $f$ then performs model aggregation following algorithms such as FedAvg (McMahan et al., 2017) to update the global model $\omega_{f}(t)$. Uploads received beyond this time point will not be accepted, and corresponding participants will not receive payment. This is especially important for meeting the deadline $D_{f}$. The federation then updates participants’ reputation scores based on the marginal improvement between global model performance and participant’s local model performance. The FL training process continues until either the predefined deadline at $t=D_{f}$ is reached or until the task is completed, whichever is earlier.
@@ 3.2. Energy Model and Execution Cost Function
(9) $E^{cmp}_{f}(t)=\sum^{N}_{n=1}P(n,t)M\zeta_{f}G_{f}^{2},\forall N\in\mathcal{N}$
where $G_{f}$ represents $f$’s CPU operating frequency, and $\zeta_{f}$ denotes the effective load capacitance of the $f$’s computation chip-set.
@@ 5. Analysis
In this section, we perform theoretical analysis on the properties of GPS-AFL, including incentive compatibility (IC), individual rationality (IR), computational time complexity, and budget feasibility.
Incentive Compatibility: Participants have no incentive to misreport their true valuations under GPS-AFL. According to (Myerson, 1981), a reverse auction is incentive compatible if and only if the following two conditions are met. Firstly, its selection criteria exhibit monotonicity, which means that if $b_{n}(t)$ is the winning bid, then any bid lower than $b_{n}(t)$ must also win the bid and be selected. Secondly, if participant $n$’s bid is higher than the winning bid $b_{n}(t)>b^{win}(t)$, then $n$ will not be selected to win the auction.
@@
According to Lines 4-9 of Algorithm 1, when $Q_{f}(t)\leqslant 0$, if any participant bids higher than the winning bid (i.e., $b_{n}(t)>b^{win}(t)$), the participant is rejected, resulting in 00 utility for that participant. If any participant $n$ bids lower than the minimum bid (i.e., $b_{n}(t)<b^{win}(t)$), the participant’s bid will become the new winning bid $b_{n}(t)=b^{win}(t)$. However, by doing so, $n$ will receive lower reward than previously possible, resulting in a sub-optimal $U_{n}(t)$. Hence, it is optimal for all participants to bid truthfully. ∎
@@ 6.2. Results and Discussion
Table 4 presents the experimental result concerning a more difficult FL task, based on the EMNIST balanced dataset. The number of participants available has been reduced to $20$. To save on communication costs, each participant will train for $3$ local epochs, with batch size $64$. Consequently, $t_{t-r}$ tripled to become $0.3$. The remaining parameters remained unaltered. The FL model deployed for this FL task is identical to the one used in the FMNIST FL task. It is worth noting that our approach can be extended to other FL tasks by leveraging on alternative available benchmarks like the open-sourced benchmarking framework offered by (Zeng et al., 2023). As shown in Table 4, it remains evident that GPS-AFL maintains its superiority over other state-of-the-art methodologies while keeping costs low. GPS-AFL was able to attain a total utility of $4.44\%$, $30.12\%$, $74.31\%$, $63.23\%$ and $43.27\%$ more points higher than ‘RRAFL’, ‘ROBO-AFL’, ‘OORT’, ‘SV’ and ‘Greedy’ approaches, respectively. While simultaneously incurring $36.38\%$, $57.18\%$, $310.98\%$, $399.31\%$ and $351.30\%$ lesser costs than the other approaches in the same order, respectively.

### v2

@@ 3. System Model of GPS-AFL
Step 4: FL Model Training. The optimization of the global loss function is achieved by minimizing the weighted mean of the participants’ local loss functions (McMahan et al., 2017), denoted as $\ell_{n}(\omega_{n}(t))$. As such, each selected participant train its local model, $\omega_{n}(t)$, for a single time step using its own local dataset and compute resources, in attempt to minimize $\ell_{n}(\omega_{n}(t))$ within the allocated time. The actual time allocated for training will be shorter than the actual time step $t$ as time has to be allocated for participants to upload their updated model weights. We use $t_{r}$, where $0<r\leqslant t$, to denote the actual time period used for model training and $t-r$ would be the amount of communication time for participants to upload their model weights.
Step 5: FL Model Aggregation. Participants are required to submit their local model updates to the FL server $f$ at the end of $t_{r}$. $f$ then performs model aggregation following algorithms such as FedAvg (McMahan et al., 2017) to update the global model $\omega_{f}(t)$. Uploads received beyond this time point will not be accepted, and corresponding participants will not receive payment. This is especially important for meeting the deadline $D_{f}$. The federation then updates participants’ reputation scores based on the marginal improvement between global model performance and participant’s local model performance. The FL training process continues until either the predefined deadline at $t=D_{f}$ is reached or until the task is completed, whichever is earlier.
@@ 3.2. Energy Model and Execution Cost Function
(9) $E^{cmp}_{f}(t)=\sum^{N}_{n=1}P(n,t)M_{n}\varsigma\zeta_{f}G_{f}^{2},\forall n\in\mathcal{N}$
where $G_{f}$ represents $f$’s CPU operating frequency and $\varsigma$ denotes the number of CPU cycles required for FL server to perform the model aggregation for each DO’s model update. $M_{n}$ represents the size of model update provided by DO $n$. Finally, $\zeta_{f}$ denotes the effective load capacitance of the $f$’s computation chip-set.
@@ 5. Analysis
In this section, we perform theoretical analysis on the properties of GPS-AFL, truthfulness, individual rationality (IR), computational time complexity, and budget feasibility.
Truthfulness: Participants would have no incentive to report a bid other than their true private valuations under GPS-AFL.
@@
According to Lines 4-9 of Algorithm 1, suppose $Q_{f}(t)\leqslant 0$ and that there is sufficient budget. Assume participant $n$ placed the lowest bid based on its private valuation and wins the round (i.e., $b_{n}(t)=b^{win}(t)$). If $n$ bids any higher than $b_{n}$ then $n$ will be rejected resulting in $P(n,t)=0$ and consequently 00 utility for that participant. Conversely, if $n$ places a lower bid (i.e., $b^{\prime}_{n}(t)<b_{n}(t)$ where $b^{\prime}_{n}(t)$ is the new bid) and wins, then $n$ will receive lower reward than previously possible, resulting in a sub-optimal or negative $U_{n}(t)$. While $b^{\prime}_{n}(t)>b_{n}(t)$ and $b^{\prime}_{n}(t)=b^{win}(t)$ might occur, $n$ would not know $b^{win}(t)$ before the round ends, preventing such bid manipulation. Moreover, due to nature of reverse-AFL, the bid price will eventually decrease. Hence, there is no incentive for any rational participant to not bid truthfully. ∎
@@ 6.2. Results and Discussion
Table 4 presents the experimental result concerning a more difficult FL task, based on the EMNIST balanced dataset. The number of participants available has been reduced to $20$. To save on communication costs, each participant will train for $3$ local epochs, with batch size $64$. Consequently, $t_{r}$ tripled to become $3$. The remaining parameters remained unaltered. The FL model deployed for this FL task is identical to the one used in the FMNIST FL task. It is worth noting that our approach can be extended to other FL tasks by leveraging on alternative available benchmarks like the open-sourced benchmarking framework offered by (Zeng et al., 2023). As shown in Table 4, it remains evident that GPS-AFL maintains its superiority over other state-of-the-art methodologies while keeping costs low. GPS-AFL was able to attain a total utility of $4.44\%$, $30.12\%$, $74.31\%$, $63.23\%$ and $43.27\%$ more points higher than ‘RRAFL’, ‘ROBO-AFL’, ‘OORT’, ‘SV’ and ‘Greedy’ approaches, respectively. While simultaneously incurring $36.38\%$, $57.18\%$, $310.98\%$, $399.31\%$ and $351.30\%$ lesser costs than the other approaches in the same order, respectively.